# Large Language Models Develop Novel Social Biases Through Adaptive Exploration

### v2

@@ Abstract
As large language models (LLMs) are adopted into frameworks that grant them the capacity to make real decisions, it is increasingly important to ensure that they are unbiased. In this paper, we argue that the predominant approach of simply removing existing biases from models is not enough. Using a paradigm from the psychology literature, we demonstrate that LLMs can spontaneously develop novel social biases about artificial demographic groups even when no inherent differences exist. These biases result in highly stratified task allocations, which are less fair than assignments by human participants and are exacerbated by newer and larger models. In social science, emergent biases like these have been shown to result from exploration-exploitation trade-offs, where the decision-maker explores too little, allowing early observations to strongly influence impressions about entire demographic groups. To alleviate this effect, we examine a series of interventions targeting model inputs, problem structure, and explicit steering. We find that explicitly incentivizing exploration most robustly reduces stratification, highlighting the need for better multifaceted objectives to mitigate bias. These results reveal that LLMs are not merely passive mirrors of human social biases, but can actively create new ones from experience, raising urgent questions about how these systems will shape societies over time.
@@ 1 Introduction
[…]
@@
In this paper, we argue that removing existing biases is only one aspect of the problem. Like people, LLMs can also invent novel biases that influence human and agent behavior. Stereotype biases in humans can naturally emerge through experiences that constrain exploration (bai_globally_nodate; bai_costly_2025; fang2011theories; merton1948self; schelling_dynamic_1971): residents search only familiar neighborhoods, reinforcing segregation (krysan2017cycle); police repeatedly patrol high-crime areas, disproportionately arresting minorities (lum2016predict); managers avoid hiring unconventional candidates, maintaining incorrect beliefs (baek2023feedback); and individuals view a group negatively after one bad encounter, escalating conflicts (denrell_adaptation_2001). This mechanism parallels the exploration-exploitation dilemma in reinforcement learning (ensign_runaway_2018; sutton1998reinforcement): when iteratively facing choices with multiple options, each choice is costly but informative, forcing decision-makers to balance exploring novel options with exploiting what worked before. This phenomena becomes pertinent at a time when foundation models are being integrated into agentic frameworks, letting them retain persistent belief states across interactions, while also granting them autonomy to make decisions with limited human oversight (krishnamurthy_can_2024; laskin2022context; raparthy2023generalization; shinn2023reflexion).
@@
We illustrate this process of developing novel biases using a hiring game paradigm from the psychology literature (bai_globally_nodate; bai_costly_2025). Participants act as hiring managers to allocate a series of jobs, each of which has candidates from four artificial demographic groups, and they are rewarded for how many hired candidates succeed. Jobs are split into four types along two psychological dimensions, warmth and competence (fiske2002model), following human data from Bai et al. (2025). For example, doctors are seen as trustworthy and competent while janitors are viewed as less so (koenig2014evidence). Unknown to the participant, all candidates are equally likely to succeed with probability at each job. However, as participants explore by assigning candidates to roles and receive feedback on whether they succeed, these early observations often lead them to form inaccurate impressions about the underlying traits of each group, leading them to stratify candidates by assigning different groups to different job types. In other words, people do not explore enough to remove biases caused by inherently random feedback, causing them to treat groups unequally despite no real differences. Afterwards, people retained these biases, rating certain groups as more competent or caring than others. This process demonstrates how humans can develop new biases simply from engaging in sequential decision-making with noisy outcomes.
@@
Our findings reflect a general, recurring theme in optimization and AI — that stronger optimizers require better-formulated goals (amodei2016concrete; hadfield2017inverse; manheim2018categorizing; pan2022effects; smith2006optimizer). As a concrete example, consider the contrast between newspapers and social media, which share the objective of increasing audience engagement. While newspapers were limited by lack of feedback, social media platforms used closed-loop optimization with user data to improve recommendations—but this led to negative societal consequences such as echo chambers and polarization (allcott2020welfare; bakshy2015exposure; cinelli2021echo). Our results show that LLMs as optimizers have also outgrown simple reasoning objectives. To adapt to the improved capabilities that state-of-the-art models provide, we believe that holistic objectives that incorporate societal values (bai2022constitutional; klingefjord2024human) are imperative to ensure that AI systems stay unbiased as they explore and interact with the world.
@@ 2.1 Quantifying and addressing biases in LLMs
[…]
@@ 2.2 Challenges for exploration with LLMs
In-context learning illustrates how LLMs can generalize from very few examples without training, leading to superior performance on many tasks (akyürek2023what; brown2020language; shi2024why). However, in this paradigm, LLMs have also displayed notable shortcomings when operating in unfamiliar distributions or on tasks that require generalization beyond surface patterns. For example, in multi-armed bandit tasks, LLMs tend to fixate on the same option that first results in a successful reward, even though this is suboptimal (krishnamurthy_can_2024; pan_large_2025; schmied2025llmsgreedyagentseffects). LLMs can also make spurious and incorrect generalizations from confounded in-context data, prioritizing surface-level features such as sentiment (fei_mitigating_2023), length (schoch_-context_2025), or those favored in its priors (si_measuring_2023). More broadly, LLMs display inductive biases toward simpler or more common patterns (mccoy_embers_2024; mccoy2024languagemodeloptimizedreasoning), which can lead to them over-indexing on such patterns within in-context data (li_patterns_2025; liu2025mind). Together, these results highlight how limited exploration—through fixation, spurious correlations, or early lock-in on presumed patterns—remains a central bottleneck to robust generalization.
@@ 3.1 Hiring game paradigm
This simple contextual multi-armed bandit setup from Bai et al. (2025) is designed to strip away existing biases: participants belonged to none of the groups—reducing in-group loyalty (brewer1979group), clear instructions and short trials minimized cognitive load (macrae1994stereotypes), and job candidates had equal population sizes to prevent data imbalance (fiedler2000beware). Crucially, unknown to participants, the odds of success were identical for every group and every job. At each round, whether any job is a good fit for any selected applicant is a random variable sampled from .
In the original experiment, human participants failed to realize that there were no meaningful differences among groups. Instead, they became entrenched in their own successes: once they observed that a Tufa was a good doctor or a Weki worked well as a janitor, participants kept repeating similar choices rather than exploring alternatives. In doing so, they inadvertently built a stratified city of their own making, and created new mental stereotypes imagining Tufas as warm and competent while casting Wekis as untrustworthy and incompetent (bai_costly_2025). This experiment provides the baseline human data for our evaluation of LLMs (see Appendix C for details), which we test using the same hiring task.
@@ 3.2 Metrics
Throughout this section, let denote the set of demographic groups, the collection of independent runs of the hiring game, and the set of 4 job classes: high competence and high warmth (e.g., doctor), high competence and low warmth (e.g., lawyer), low competence and high warmth (e.g., childcare aide), and low competence and low warmth (e.g., janitor) (bai_costly_2025; fiske2002model; fiske2014gaining; koenig2014evidence). For each group in run , we write for its empirical allocation distribution over the job classes, and for the uniform distribution on . and JSD denote entropy and Jensen-Shannon divergence over probability distributions, respectively, with all logarithms calculated using base 2.
@@ 4.1 Models and hyperparameters
We examined a variety of state-of-the-art LLMs and their predecessors, both proprietary and open-source: GPT-[3.5, 4o], Claude [3 Haiku, 4 Sonnet], Gemini [1.5, 2.0, 2.5] Flash, Qwen 2.5-[7B, 72B] Instruct Turbo, Llama [3.2 3B, 11B, 90B, 4 Scout 17B-16E, 4 Maverick 17B-128E] (frontier models of each family are in bold). In addition, we tested two reasoning models, one proprietary—OpenAI o3, and one open-source—DeepSeek-R1. Each model was prompted at its default temperature, with both direct and chain-of-thought prompting (CoT; wei_chain--thought_2022). For reasoning models, the default medium reasoning effort was used. For each model and prompt type, we collected runs of the 40-round hiring game from Section 3.1, with the order of jobs shuffled each run. Prompts are in Appendix A.1.
@@ 4.2 Results
In experiments across each model family {Claude, GPT, Gemini, Llama3.2, Llama4, Qwen2.5}, we observe that newer and larger models stratified statistically significantly more as measured by both SI and BGD (Figure 3). For instance, Claude 4 Sonnet’s SI was more than eight times Claude 3 Haiku’s in the direct prompting condition. This runs contrary to results on single-prompt bias benchmarks such as BBQ, where newer and larger models consistently demonstrate higher performance than predecessors (noauthor_safety_nodate; liang_holistic_2023; parrish_bbq_2022). We provide a visualization of the strong negative correlation between BBQ score and negative SI (adjusted so that higher = better) in Figure 4. We also provide a visualization for how newer models within each family stratify more using single-run rank-ordered job allocation matrices (Appendix D). Our finding that stronger models are more at risk of developing emergent biases highlights the need to specifically attend to this new kind of bias.
@@ 5.1 System-level interventions
CoT has shown promise in encouraging exploration and reducing bias (10.1145/3715275.3732208; krishnamurthy_can_2024), and is a general strategy to improve performance (wei_chain--thought_2022). While CoT decreased stratification in most frontier models (Figure 2), these changes were often not statistically significant. With CoT, Qwen 2.5 72B—the lowest SI frontier model—reduced stratification to within human ranges. However, all outcomes were still far more stratified than fair random assignments.
@@
Another standard strategy to encourage randomness is to increase model temperature (du2025optimizing). We test this by prompting each frontier model (except Claude 4 Sonnet whose maximal temperature is ) with an increased temperature of for runs. We report only direct prompting results, as CoT devolved outputs into gibberish after 7-10 rounds at and for all models. For direct prompts, increasing the temperature to did not produce statistically significant reductions in stratification for Gemini 2.5 (), GPT-4o (), or Llama 4 Maverick (). While we observed a statistically significant decrease in stratification for Qwen 2.5-72B (), the resultant SI of remained above the human baseline—well within the high-stratification regime.
@@ 5.2 Structural interventions
To confirm that stratification is not caused by our specific setup, we test two additional settings with similar multi-turn decisions: refugee resettlement (bansak_improving_2018; bansak_how_2016) and military conscript assignment (sorlie2020person). Starting from the same multi-turn allocation paradigm, we replaced categorized jobs with either geographically-clustered cities in a country or military camps from different divisions. In the resettlement setting, we also replaced the fictional demographics with low-resource indigenous ethnicities from Central Asia for further realism, confirming that initial biases across ethnicities are spurious (across all conditions GASI ). While the original experiment from bai_costly_2025 used a points system for successful job assignments to incentivize participants, these incentives are not necessary for LLMs. Thus, our new settings remove the points system and only instruct the LLM to maximize successful assignments. See Appendices A.4 and A.5 for prompts and Appendix E for full results.
@@
Another case to consider is scenarios where the LLM has access to richer information beyond group labels alone. Real-world decision making can involve multiple dimensions of context, and incorporating additional features allows us to explore if stratification arises when models can explain observations using other available features. We examined this question using the refugee resettlement setting with established realistic feature from bansak_improving_2018; bansak_how_2016: age and education. For experiment details and prompts, see Appendix A.4.
@@
While we use the most prevalent features (age, education) for the resettlement task in our previous analysis (bansak_improving_2018; bansak_how_2016), in real world applications a myriad of features could be available for individuals. Thus, it is imperative to distinguish whether arbitrary features equally increase exploration by expanding the hypothesis space, or if LLMs selectively adjust stratification based on the additional features’ contextual importance. To examine this, we replicate the resettlement experiment using two comparatively less salient features: hair color and tattoo shape (martin2014spontaneous). We observe substantially higher levels of stratification with these features (Figure 6(b)), with mean reductions in SI of 0.25, 0.44, and 0.42 for hair color, tattoo shape, and both, compared to 0.43, 0.59, and 0.70 for age, education, and both. This suggests that LLMs are sensitive to the contextual importance of additional features when determining allocations, meaning that in real applications, reductions in stratification are conditioned on the quality of known features in available data.

### v3

@@ Abstract
As large language models (LLMs) are adopted into frameworks that grant them the capacity to make real decisions, it is increasingly important to ensure that they are unbiased. In this paper, we argue that the predominant approach of simply removing existing biases from models is not enough. Using a paradigm from the psychology literature, we demonstrate that LLMs can spontaneously develop novel social biases about artificial demographic groups even when no inherent differences exist. These biases result in highly stratified task allocations, which are less fair than assignments by human participants and are exacerbated by newer and larger models. In social science, emergent biases like these have been shown to result from exploration-exploitation trade-offs, where the decision-maker explores too little, allowing early observations to strongly influence impressions about entire demographic groups. To alleviate this effect, we examine a series of interventions targeting model inputs, problem structure, and explicit steering. We find that explicitly incentivizing exploration most robustly reduces stratification, highlighting the need for better multifaceted objectives to mitigate bias. These results reveal that LLMs are not merely passive mirrors of human social biases, but can actively create new ones from experience, raising urgent questions about how these systems will shape societies over time. 00footnotetext: Our code is available at github.com/addisonwu05/LLM-Emergent-Stereotypes.
@@ 1 Introduction
As LLMs become embedded in everyday applications across countless tasks, it is imperative for them to be unbiased, meaning that they treat people equally across racial, gender, and other social groups. This is critical because biased behavior in such systems can perpetuate and amplify existing societal inequities, undermine user trust, and lead to systematically unequal access to resources and opportunities. However, current LLMs are biased: they mirror existing human biases (e.g., Bolukbasi et al., 2016; Caliskan et al., 2017; Dhamala et al., 2021; Nadeem et al., 2021; Tamkin et al., 2023), and many efforts have been dedicated towards removing these biases (e.g., Bordia and Bowman, 2019; Guo et al., 2022; Liang et al., 2021; Meade et al., 2022; Yu et al., 2023). This process has proven to be challenging, as models that pass benchmarks continue to reveal subtle discriminatory behaviors (Bai et al., 2025b; Hofmann et al., 2024; Ji et al., 2025; Zipperling et al., 2025).
@@
In this paper, we argue that removing existing biases is only one aspect of the problem. Like people, LLMs can also invent novel biases that influence human and agent behavior. Stereotype biases in humans can naturally emerge through experiences that constrain exploration (Bai et al., 2022a; 2025a; Fang and Moro, 2011; Merton, 1948; Schelling, 1971): residents search only familiar neighborhoods, reinforcing segregation (Krysan and Crowder, 2017); police repeatedly patrol high-crime areas, disproportionately arresting minorities (Lum and Isaac, 2016); managers avoid hiring unconventional candidates, maintaining incorrect beliefs (Baek and Makhdoumi, 2023); and individuals view a group negatively after one bad encounter, escalating conflicts (Denrell and March, 2001). This mechanism parallels the exploration-exploitation dilemma in reinforcement learning (Ensign et al., 2018; Sutton et al., 1998): when iteratively facing choices with multiple options, each choice is costly but informative, forcing decision-makers to balance exploring novel options with exploiting what worked before. This phenomena becomes pertinent at a time when foundation models are being integrated into agentic frameworks, letting them retain persistent belief states across interactions, while also granting them autonomy to make decisions with limited human oversight (Krishnamurthy et al., 2024; Laskin et al., 2023; Raparthy et al., 2024; Shinn et al., 2023).
@@
We illustrate this process of developing novel biases using a hiring game paradigm from the psychology literature (Bai et al., 2022a; 2025a). Participants act as hiring managers to allocate a series of jobs, each of which has candidates from four artificial demographic groups, and they are rewarded for how many hired candidates succeed. Jobs are split into four types along two psychological dimensions, warmth and competence (Fiske et al., 2002), following human data from Bai et al. (2025). For example, doctors are seen as trustworthy and competent while janitors are viewed as less so (Koenig and Eagly, 2014). Unknown to the participant, all candidates are equally likely to succeed with probability at each job. However, as participants explore by assigning candidates to roles and receive feedback on whether they succeed, these early observations often lead them to form inaccurate impressions about the underlying traits of each group, leading them to stratify candidates by assigning different groups to different job types. In other words, people do not explore enough to remove biases caused by inherently random feedback, causing them to treat groups unequally despite no real differences. Afterwards, people retained these biases, rating certain groups as more competent or caring than others. This process demonstrates how humans can develop new biases simply from engaging in sequential decision-making with noisy outcomes.
@@
Our findings reflect a general, recurring theme in optimization and AI — that stronger optimizers require better-formulated goals (Amodei et al., 2016; Hadfield-Menell et al., 2017; Manheim and Garrabrant, 2018; Pan et al., 2022; Smith and Winkler, 2006). As a concrete example, consider the contrast between newspapers and social media, which share the objective of increasing audience engagement. While newspapers were limited by lack of feedback, social media platforms used closed-loop optimization with user data to improve recommendations—but this led to negative societal consequences such as echo chambers and polarization (Allcott et al., 2020; Bakshy et al., 2015; Cinelli et al., 2021). Our results show that LLMs as optimizers have also outgrown simple reasoning objectives. To adapt to the improved capabilities that state-of-the-art models provide, we believe that holistic objectives that incorporate societal values (Bai et al., 2022c; Klingefjord et al., 2024) are imperative to ensure that AI systems stay unbiased as they explore and interact with the world.
@@ 2.1 Quantifying and addressing biases in LLMs
Stereotype biases in language models are well recognized as a long-standing problem, from word embeddings (Bolukbasi et al., 2016; Caliskan et al., 2017) to autoregressive models (Dhamala et al., 2021; Liang et al., 2023; Nadeem et al., 2021; Huang et al., 2025). To evaluate these biases, benchmarks have mainly focused on existing categories embedded in society, such as race (Hofmann et al., 2024; Wang et al., 2023), gender and sexual orientation (Ovalle et al., 2023; Wan et al., 2023), age (Tamkin et al., 2023), religion (Abid et al., 2021), occupation (Kirk et al., 2021), and cultural background (Shen et al., 2024). To reduce these biases, intervention techniques also target known stereotypes by creating alignment datasets (Bai et al., 2022b; Zhang et al., 2025), editing model activations (Prakash and Roy, 2024; Sun et al., 2025; Yu and Ananiadou, 2025), or prompting (Si et al., 2023). While useful for addressing existing biases, these approaches cannot capture or address new forms of bias that emerge as models interact with the world and adapt their beliefs. Here, we show that LLMs can generate entirely novel and potentially problematic biases, unseen in any data.
@@ 2.2 Challenges for exploration with LLMs
In-context learning illustrates how LLMs can generalize from very few examples without training, leading to superior performance on many tasks (akyürek2023what; Brown et al., 2020; Shi et al., 2024). However, in this paradigm, LLMs have also displayed notable shortcomings when operating in unfamiliar distributions or on tasks that require generalization beyond surface patterns. For example, in multi-armed bandit tasks, LLMs tend to fixate on the same option that first results in a successful reward, even though this is suboptimal (Krishnamurthy et al., 2024; Pan et al., 2025; Schmied et al., 2025). LLMs can also make spurious and incorrect generalizations from confounded in-context data, prioritizing surface-level features such as sentiment (Fei et al., 2023), length (Schoch and Ji, 2025), or those favored in its priors (Si et al., 2023). More broadly, LLMs display inductive biases toward simpler or more common patterns (McCoy et al., 2024a; b), which can lead to them over-indexing on such patterns within in-context data (Li et al., 2025; Liu et al., 2025). Together, these results highlight how limited exploration—through fixation, spurious correlations, or early lock-in on presumed patterns—remains a central bottleneck to robust generalization.
@@ 3.1 Hiring game paradigm
This simple contextual multi-armed bandit setup from Bai et al. (2025) is designed to strip away existing biases: participants belonged to none of the groups—reducing in-group loyalty (Brewer, 1979), clear instructions and short trials minimized cognitive load (Macrae et al., 1994), and job candidates had equal population sizes to prevent data imbalance (Fiedler, 2000). Crucially, unknown to participants, the odds of success were identical for every group and every job. At each round, whether any job is a good fit for any selected applicant is a random variable sampled from .
In the original experiment, human participants failed to realize that there were no meaningful differences among groups. Instead, they became entrenched in their own successes: once they observed that a Tufa was a good doctor or a Weki worked well as a janitor, participants kept repeating similar choices rather than exploring alternatives. In doing so, they inadvertently built a stratified city of their own making, and created new mental stereotypes imagining Tufas as warm and competent while casting Wekis as untrustworthy and incompetent (Bai et al., 2025a). This experiment provides the baseline human data for our evaluation of LLMs (see Appendix C for details), which we test using the same hiring task.
@@ 3.2 Metrics
Throughout this section, let denote the set of demographic groups, the collection of independent runs of the hiring game, and the set of 4 job classes: high competence and high warmth (e.g., doctor), high competence and low warmth (e.g., lawyer), low competence and high warmth (e.g., childcare aide), and low competence and low warmth (e.g., janitor) (Bai et al., 2025a; Fiske et al., 2002; Fiske and Dupree, 2014; Koenig and Eagly, 2014). For each group in run , we write for its empirical allocation distribution over the job classes, and for the uniform distribution on . and JSD denote entropy and Jensen-Shannon divergence over probability distributions, respectively, with all logarithms calculated using base 2.
@@ 4.1 Models and hyperparameters
We examined a variety of state-of-the-art LLMs and their predecessors, both proprietary and open-source: GPT-[3.5, 4o], Claude [3 Haiku, 4 Sonnet], Gemini [1.5, 2.0, 2.5] Flash, Qwen 2.5-[7B, 72B] Instruct Turbo, Llama [3.2 3B, 11B, 90B, 4 Scout 17B-16E, 4 Maverick 17B-128E] (frontier models of each family are in bold). In addition, we tested two reasoning models, one proprietary—OpenAI o3, and one open-source—DeepSeek-R1. Each model was prompted at its default temperature, with both direct and chain-of-thought prompting (CoT; Wei et al., 2022). For reasoning models, the default medium reasoning effort was used. For each model and prompt type, we collected runs of the 40-round hiring game from Section 3.1, with the order of jobs shuffled each run. Prompts are in Appendix A.1.
@@ 4.2 Results
In experiments across each model family {Claude, GPT, Gemini, Llama3.2, Llama4, Qwen2.5}, we observe that newer and larger models stratified statistically significantly more as measured by both SI and BGD (Figure 3). For instance, Claude 4 Sonnet’s SI was more than eight times Claude 3 Haiku’s in the direct prompting condition. This runs contrary to results on single-prompt bias benchmarks such as BBQ, where newer and larger models consistently demonstrate higher performance than predecessors (Center for Research on Foundation Models, 2025; Liang et al., 2023; Parrish et al., 2022). We provide a visualization of the strong negative correlation between BBQ score and negative SI (adjusted so that higher = better) in Figure 4. We also provide a visualization for how newer models within each family stratify more using single-run rank-ordered job allocation matrices (Appendix D). Our finding that stronger models are more at risk of developing emergent biases highlights the need to specifically attend to this new kind of bias.
@@ 5.1 System-level interventions
CoT has shown promise in encouraging exploration and reducing bias (Gupta et al., 2025; Krishnamurthy et al., 2024), and is a general strategy to improve performance (Wei et al., 2022). While CoT decreased stratification in most frontier models (Figure 2), these changes were often not statistically significant. With CoT, Qwen 2.5 72B—the lowest SI frontier model—reduced stratification to within human ranges. However, all outcomes were still far more stratified than fair random assignments.
@@
Another standard strategy to encourage randomness is to increase model temperature (Du et al., 2025). We test this by prompting each frontier model (except Claude 4 Sonnet whose maximal temperature is ) with an increased temperature of for runs. We report only direct prompting results, as CoT devolved outputs into gibberish after 7-10 rounds at and for all models. For direct prompts, increasing the temperature to did not produce statistically significant reductions in stratification for Gemini 2.5 (), GPT-4o (), or Llama 4 Maverick (). While we observed a statistically significant decrease in stratification for Qwen 2.5-72B (), the resultant SI of remained above the human baseline—well within the high-stratification regime.
@@ 5.2 Structural interventions
To confirm that stratification is not caused by our specific setup, we test two additional settings with similar multi-turn decisions: refugee resettlement (Bansak et al., 2018; 2016) and military conscript assignment (Sørlie et al., 2020). Starting from the same multi-turn allocation paradigm, we replaced categorized jobs with either geographically-clustered cities in a country or military camps from different divisions. In the resettlement setting, we also replaced the fictional demographics with low-resource indigenous ethnicities from Central Asia for further realism, confirming that initial biases across ethnicities are spurious (across all conditions GASI ). While the original experiment from Bai et al. (2025a) used a points system for successful job assignments to incentivize participants, these incentives are not necessary for LLMs. Thus, our new settings remove the points system and only instruct the LLM to maximize successful assignments. See Appendices A.4 and A.5 for prompts and Appendix E for full results.
@@
Another case to consider is scenarios where the LLM has access to richer information beyond group labels alone. Real-world decision making can involve multiple dimensions of context, and incorporating additional features allows us to explore if stratification arises when models can explain observations using other available features. We examined this question using the refugee resettlement setting with established realistic feature from Bansak et al. (2018; 2016): age and education. For experiment details and prompts, see Appendix A.4.
@@
While we use the most prevalent features (age, education) for the resettlement task in our previous analysis (Bansak et al., 2018; 2016), in real world applications a myriad of features could be available for individuals. Thus, it is imperative to distinguish whether arbitrary features equally increase exploration by expanding the hypothesis space, or if LLMs selectively adjust stratification based on the additional features’ contextual importance. To examine this, we replicate the resettlement experiment using two comparatively less salient features: hair color and tattoo shape (Martin et al., 2014). We observe substantially higher levels of stratification with these features (Figure 6(b)), with mean reductions in SI of 0.25, 0.44, and 0.42 for hair color, tattoo shape, and both, compared to 0.43, 0.59, and 0.70 for age, education, and both. This suggests that LLMs are sensitive to the contextual importance of additional features when determining allocations, meaning that in real applications, reductions in stratification are conditioned on the quality of known features in available data.